An object which is released from a hot-air balloon at a height experiences a force of air resistance which is directly proportional to the velocity, v , after t seconds satisfies the differential equation dv/dt + 0.25v = 10, where v is in meter per second and t is seconds. Assuming that the initial velocity of the object is 0, find v in terms of t.
1
Expert's answer
2020-10-26T20:10:16-0400

"\displaystyle

\frac{\mathrm{d}v}{\mathrm{d}t} + 0.25v = 10\\

\textsf{This can be solved by multiplying}\\\textsf{both sides of the differential equation}\\\textsf{by an integrating factor.}\\


\frac{\mathrm{d}v}{\mathrm{d}t} + 0.25v = 10\\


I.F = \textsf{Integrating factor}\\



I.F = e^{\int 0.25 \, \mathrm{d}t} = e^{\frac{t}{4}}\\

\textsf{The solution to the first order linear}\\\textsf{differential equation is thus;}\\


v\, I.F = \int 10 I.F \, \mathrm{d}t\\

\begin{aligned}
ve^{\frac{t}{4}} &= \int 10\,e^{\frac{t}{4}} \, \mathrm{d}t\\
ve^{\frac{t}{4}} &= 10 \times 4 e^{\frac{t}{4}} + C\\
ve^{\frac{t}{4}} &= 40e^{\frac{t}{4}} + C\\
v&= 40 + Ce^{-\frac{t}{4}}\\
v(t) &= 40 + Ce^{-\frac{t}{4}}
\end{aligned}\\


\textsf{It is given that} \,v(0) = 0\\


v(0)= 40 + C = 0 \implies C = -40\\


\therefore v(t)= 40 - 40e^{-\frac{t}{4}} = 40\left(1 - e^{-\frac{t}{4}}\right)"